THE Ã − X̃ AND B̃ − X̃ ABSORPTIONS OF NO_{3} TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

Absorptions arising from the Ã − X̃ transition of normal and isotopically substituted NO_{3} have been observed between 7500 and 9500 cm^{−1}. Details of the spectra will be discussed and assignments will be proposed. Absorptions arising from the B̃ − X̃ transition of NO_{3}, with band origin near 15 000 cm^{−1}, have also been observed for the normal species and two of its isotopologues which possess D_{3h} symmetry. As in the gas phase, the absorptions are broadened because of predissociation. The observed band structure corresponds closely with that reported for the gas-phase molecule.